OpenStudy (anonymous):

express the ratio 45 seconds to 30 minutes as a fraction in simplest form

OpenStudy (anonymous):

1:5?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/51471effe4b04cdfc582d507 There you go

OpenStudy (anonymous):

\[\frac{ 45 }{ 1800}\div \frac{ 45 }{ 45 }=\frac{ 45\div45 }{ 1800\div45 }=\frac{ 1 }{ 40}\]